# On tori periods of Weil representations of unitary groups

### v1

@@ 1. Introduction
For every choice of $\mathbf{E}$, $\alpha$, and $\beta$, as above, there exists a non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and an admissible embedding $i\colon L_{\mathbf{E}}^{1}\to\operatorname{U}(\mathrm{H})$, such that $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is not identically zero on $\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$, if and only if the following conditions hold:
@@
    Character compatibility: $\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{\mathbb{A}_{\mathbf{F}}}^{-1}\cdot\alpha_{\restriction_{L_{\mathbf{F}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)}}.$
@@
    $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\chi_{\mathrm{W}}^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)\neq 0.$
@@
$\mathcal{V}_{\mathrm{W},\mathbf{E}}\left(\beta\right)=\bigcup_{\mathrm{H}\in\mathrm{Her}_{n}}\left\{\left(\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)\mid i\in\Sigma_{\mathbf{E},\mathrm{H}}\right\},$
serves as a substitute for the global Vogan packet appearing in the global Gan–Gross-Prasad conjectures, consisting of automorphic representations of the group and its pure inner forms. The results of this paper, combined with the exceptional theta correspondence of Bakić and Savin [3] for $\left(\mathrm{PU}_{3}\rtimes\mathbb{Z}/2\mathbb{Z}\right)\times G_{2}$, are used in order to construct local and global Arthur packets for the exceptional group $G_{2}$, see [1] and [2]. The paper is organized as follows. In Section 2, we recall the notion of $\epsilon$-hermitian spaces and the invariants attached to them. Then we discuss the classification of maximal tori in unitary groups associated to $\epsilon$-hermitian spaces, and discuss the notion of an admissible embedding. In Section 3, we recall the theta correspondence for unitary groups over local fields. We use it to define the notion of a big theta lift for characters of unitary groups of one-dimensional $\epsilon$-hermitian spaces over an étale algebra. We then explain a seesaw identity for this theta lift. This is a key ingredient needed for our main results. We then discuss the well-known results of theta lifting from $\operatorname{U}(1)$ to $\operatorname{U}(1)$, and use them to obtain similar results for theta lifting from $\operatorname{U}(1)$ to $\operatorname{U}(1)$ for a one-dimensional $\epsilon$-hermitian space over an étale algebra. In Section 4 we discuss the global analogs of the statements in Section 3. In Section 5, we state and prove our main theorems regarding toric periods of Weil representations of unitary groups. In the appendix, we prove statements regarding two embeddings of a maximal torus that are conjugate.
@@ 2.1. $\epsilon$-hermitian spaces and their corresponding unitary groups
Let $F$ be a field. We will always assume that $F$ has characteristic $\neq 2$. Let $K/F$ be a quadratic étale algebra with an involution, denoted $x\mapsto x^{c}$, whose fixed points are $F$. That is, $K$ is either a quadratic field extension of $F$, in which case the involution is the nontrivial Galois action on $K$, or $K=F\times F$, in which case the involution is given by $(x,y)\mapsto(y,x)$ for $x,y\in F$. The latter case is known as the split case, and the unitary groups for hermitian spaces over these two quadratic étale algebras will be treated separately. For $\epsilon\in\left\{\pm 1\right\}$, a finite dimensional $\epsilon$-hermitian space is a finite dimensional vector space $\mathrm{V}$ over $K$, equipped with an $\epsilon$-hermitian form, i.e., there exists a function $\left\langle\cdot,\cdot\right\rangle\colon\mathrm{V}\times\mathrm{V}\rightarrow K$, such that for all $w,x,y\in\mathrm{V}$, and $\alpha\in K$:
@@
If $\epsilon=1$, we call $\mathrm{V}$ hermitian, and if $\epsilon=-1$, we call $\mathrm{V}$ skew-hermitian. If $\delta\in K^{\times}$ is a trace zero element, that is, $\operatorname{tr}_{K/F}\left(\delta\right)=0$, consider the space $(\mathrm{W}^{\delta},\left\langle\cdot,\cdot\right\rangle_{\mathrm{W}^{\delta}})=(\mathrm{V},\left\langle\cdot,\cdot\right\rangle^{\delta})$ equipped with the form defined by $\left\langle x,y\right\rangle^{\delta}=\delta\left\langle x,y\right\rangle$. We have that the space $\mathrm{W}^{\delta}$ is a $-\epsilon$-hermitian space. Throughout the text, hermitian spaces will always be denoted using the symbol $\mathrm{V}$ or $\mathrm{H}$, and skew-hermitian spaces will always be denoted using the symbol $\mathrm{W}$. We say that the space $\mathrm{V}$ is non-degenerate if for every $x\in\mathrm{V}$, there exists $y\in\mathrm{V}$, such that $\left\langle x,y\right\rangle\neq 0$. In this case, if $T\colon\mathrm{V}\rightarrow\mathrm{V}$ is a linear map, then there exists a unique linear map $T^{\ast}\colon\mathrm{V}\rightarrow\mathrm{V}$, such that
@@ 2.1.1. Invariants of $\epsilon$-hermitian spaces
$\det(\left\langle\cdot,\cdot\right\rangle)_{\mathbf{e}}:=\det(\left\langle e_{i},e_{j}\right\rangle)_{ij}.$
If $\mathrm{V}$ is hermitian, then $\det\left(\left\langle\cdot,\cdot\right\rangle\right)_{\mathbf{e}}\in F^{\times}$ and it is well-known that the class of $\det(\left\langle\cdot,\cdot\right\rangle)_{\mathbf{e}}$ in $F^{\times}/N_{K/F}(K^{\times})$ is independent of the chosen basis $\mathbf{e}$. Hence, we may omit $\mathbf{e}$ from the notation and denote $\det\mathrm{V}=\det(\left\langle\cdot,\cdot\right\rangle)_{\mathbf{e}}\in F^{\times}/N_{K/F}\left(K^{\times}\right)$. Note that in the split case where $K=F\times F$, we have that $F^{\times}=N_{K/F}\left(K^{\times}\right)$, and therefore the invariant $\det\mathrm{V}$ is always trivial. We define the discriminant of $\mathrm{V}$ as
@@
Suppose now that $F$ is a local field (we allow both archimedean and non-archimedean local fields) and that $K/F$ is a quadratic field extension. Using the non-trivial quadratic character $\omega_{K/F}$ of $F^{\times}/N_{K/F}(K^{\times})$, we can encode the discriminant as a sign. For a hermitian space $\mathrm{V}$ as above we define
@@ 2.2. One-dimensional hermitian spaces with respect to an étale algebra
Let $L$ be an étale algebra of degree $n$ over $K$. We say that $\left(L,\sigma\right)$ is an étale algebra with involution if $\sigma\colon L\rightarrow L$ is an involution, such that for any $a\in K$, $\sigma(a)=a^{c}$. Given such an étale algebra with involution, we may define the notion of an $L$-hermitian (or $L$-skew-hermitian) space as in Section 2.1 by replacing $K$ with $L$ and the involution $x\mapsto x^{c}$ with the involution $\sigma$ in the definitions. It is well-known that any étale algebra with involution is of the form $L=L_{E}=K\otimes_{F}E$, where $E$ is an étale algebra of degree $n$ over $F$ and $\sigma\colon L_{E}\rightarrow L_{E}$ is defined on pure tensors by
@@
for $k\in K$ and $h\in E$. Henceforth we will often write $\sigma\left(x\right)=x^{c}$ for $x\in L_{E}$. Let $E$ be an étale algebra as above. For any $\lambda\in E^{\times}$, let $L_{E,\lambda}=(L_{E},\left\langle\cdot,\cdot\right\rangle_{L_{E,\lambda}})$ be the one-dimensional $L_{E}$-hermitian space equipped with the following form:
@@
As before, given a trace zero element $\delta\in L_{E}^{\times}$, we may define a one-dimensional $L_{E}$-skew-hermitian space $\mathrm{W}_{E,\lambda}^{\delta}=(L_{E},\left\langle\cdot,\cdot\right\rangle_{\mathrm{W}_{E,\lambda}^{\delta}})$ by setting $\left\langle x,y\right\rangle_{\mathrm{W}_{E,\lambda}^{\delta}}=\delta\left\langle x,y\right\rangle_{L_{E,\lambda}}$. Consider the unitary group $\operatorname{U}(L_{E,\lambda})$, consisting of all $L_{E}$-linear maps that preserve the $L_{E}$-hermitian structure on $L_{E,\lambda}$. It is naturally isomorphic to the group of norm one elements of $L_{E}$:
@@
Similarly, the unitary group $\operatorname{U}(\mathrm{W}_{E,\lambda}^{\delta})$ is also naturally isomorphic to $L_{E}^{1}$. The group $L_{E}^{1}$ will be fundamental for defining maximal tori in unitary groups in the next sections. If $E=\prod_{j=1}^{m}F_{j}$, where for every $j$, $F_{j}/F$ is a field extension, and $\lambda=\left(\lambda_{1},\dots,\lambda_{m}\right)\in\prod_{j=1}^{m}F_{j}^{\times}$, then we have that $\operatorname{U}(L_{E,\lambda})=\prod_{j=1}^{m}\operatorname{U}(L_{F_{j},\lambda_{j}})$, which is naturally isomorphic to $\prod_{j=1}^{m}L_{F_{j}}^{1}$. We also have a similar statement for the unitary group $\operatorname{U}(\mathrm{W}_{E,\lambda}^{\delta})$ of the skew-hermitian $L_{E}$-space $\mathrm{W}_{E,\lambda}^{\delta}$.
@@ 2.2.1. The discriminant of an étale algebra
$\det\left(\operatorname{tr}_{E/F}\left(\cdot\right)\right)_{\mathbf{e}}\coloneqq\left(\operatorname{tr}_{E/F}\left(e_{i}e_{j}\right)\right)_{i,j}.$
@@ 2.3.1. Classification of maximal tori of unitary groups
In this section, we utilize the space $\mathrm{V}_{E,\lambda}$ and the unitary group $\operatorname{U}(L_{E,\lambda})$ in order to classify the maximal tori of $\operatorname{U}(\mathrm{V})$ where $\mathrm{V}$ is a non-degenerate $n$-dimensional hermitian space. For any $x\in L_{E}$, we may consider the $K$-linear map $\mathrm{m}_{x}\colon L_{E}\rightarrow L_{E}$ defined by $\mathrm{m}_{x}\left(y\right)=xy$. For any $\lambda\in E^{\times}$ and any $x\in L_{E}^{1}$, we have that the map $\mathrm{m}_{x}$ preserves the hermitian form $\left\langle\cdot,\cdot\right\rangle_{\lambda}$. Let
@@
By [26, Proposition 2.3] we have that $\mathrm{T}$ is a maximal torus in $\operatorname{U}(\mathrm{V}_{E,\lambda})$. If $\mathrm{V}$ is a non-degenerate $n$-dimensional hermitian space, all maximal tori of $\operatorname{U}(\mathrm{V})$ can be described in a similar way. First, if $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$ is an isomorphism of hermitian spaces, then the map $\operatorname{U}(\mathrm{V}_{E,\lambda})\rightarrow\operatorname{U}(\mathrm{V})$ given by $g\mapsto r\circ g\circ r^{-1}$ is an isomorphism, and therefore
@@
is a maximal torus in $\operatorname{U}(\mathrm{V})$. For the other direction, by [26, Proposition 2.3] if $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ is a maximal torus, then there exists an $n$-dimensional étale algebra $E$ and an embedding $i^{\prime}\colon L_{E}\rightarrow\operatorname{End}\left(\mathrm{V}\right)$, such that $\mathrm{T}=i^{\prime}\left(\operatorname{Res}_{L_{E}/F}\left(\operatorname{GL}_{1}\right)\right)\cap\operatorname{U}(\mathrm{V})$, or equivalently $\mathrm{T}=i^{\prime}\left(L_{E}^{1}\right)$. Finally, by [4, Proposition 1.4.1], if $E$ is an $n$-dimensional étale algebra over $F$, then there exists an embedding $i^{\prime}\colon L_{E}\rightarrow\operatorname{End}\left(\mathrm{V}\right)$ if and only if there exists $\lambda\in E^{\times}$ and an isomorphism of hermitian spaces $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$. By the proof of [4, Proposition 1.4.1], in this case we have $i^{\prime}\left(x\right)=r\circ\mathrm{m}_{x}\circ r^{-1}$ for any $x\in L_{E}$. We summarize these results in the following theorem.
@@
Hence, the statement in the split case is equivalent to the well-known fact that maximal tori in $\operatorname{GL}_{n}\left(F\right)$ are in bijection with subgroups of the form $\left\{\left[\mathrm{m}_{x}\right]_{\mathcal{B}_{E}}\mid x\in E^{\times}\right\}$ for some étale algebra $E$ of degree $n$ over $F$ and some $F$-basis $\mathcal{B}_{E}$ of $E$, where $\left[\mathrm{m}_{x}\right]_{\mathcal{B}_{E}}$ is the matrix representing $\mathrm{m}_{x}$ with respect to the basis $\mathcal{B}_{E}$. See for example [28, Proposition 3.2.29] or [30, Section 6.1].

### v2

@@ 1. Introduction
For every choice of $\mathbf{E}$, $\alpha$, and $\beta$, as above, there exists a non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and an admissible embedding $i\colon L_{\mathbf{E}}^{1}\to\operatorname{U}(\mathrm{H})$, such that $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is not identically zero on $\Theta_{\delta,\mathrm{H},\mu,\psi}\left(\beta\right)$, if and only if the following conditions hold:
@@
    Character compatibility: $\beta=\alpha_{\restriction_{L_{\mathbf{F}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)}}.$
@@
    $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\mu^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)\neq 0.$
@@
$\mathcal{V}_{\delta,\mathbf{E}}\left(\beta\right)=\bigcup_{\mathrm{H}\in\mathrm{Her}_{n}}\left\{\left(\Theta_{\delta,\mathrm{H},\mu,\psi}\left(\beta\right),\alpha\circ i^{-1}\right)\mid i\in\Sigma_{\mathbf{E},\mathrm{H}}\right\},$
serves as a substitute for the global Vogan packet appearing in the global Gan–Gross-Prasad conjectures, consisting of automorphic representations of the group and its pure inner forms.
The results of this paper, combined with the exceptional theta correspondence of Bakić and Savin [3] for $\left(\mathrm{PU}_{3}\rtimes\mathbb{Z}/2\mathbb{Z}\right)\times G_{2}$, are used in order to construct local and global Arthur packets for the exceptional group $G_{2}$, see [1] and [2].
The paper is organized as follows. In Section 2, we recall the notion of $\epsilon$-hermitian spaces and the invariants attached to them. Then we discuss the classification of maximal tori in unitary groups associated to $\epsilon$-hermitian spaces, and discuss the notion of an admissible embedding. In Section 3, we recall the theta correspondence for unitary groups over local fields. We use it to define the notion of the big theta lift for characters of unitary groups of one-dimensional $\epsilon$-hermitian spaces over an étale algebra. We then explain a seesaw identity for this theta lift. This is a key ingredient needed for our main results. We then discuss the well-known results of theta lifting from $\operatorname{U}(1)$ to $\operatorname{U}(1)$, and use them to obtain similar results for theta lifting from $\operatorname{U}(1)$ to $\operatorname{U}(1)$ for a one-dimensional $\epsilon$-hermitian space over an étale algebra. In Section 4 we discuss the global analogs of the statements in Section 3. In Section 5, we state and prove our main theorems regarding toric periods of Weil representations of unitary groups. In the appendix, we prove statements regarding two embeddings of a maximal torus that are conjugate.
@@ 1.1. Acknowledgments
Finally, we would like to thank the anonymous referee for their thorough reading and their valuable comments and suggestions that significantly improved the mathematical exposition.
@@ 2.1. $\epsilon$-hermitian spaces and their corresponding unitary groups
Let $F$ be a field. We will always assume that $F$ has characteristic $\neq 2$. Let $K/F$ be a quadratic étale algebra with an involution, denoted $x\mapsto x^{c}$, whose fixed points are $F$. That is, $K$ is either a quadratic field extension of $F$, in which case the involution is the nontrivial Galois action on $K$, or $K=F\times F$, in which case the involution is given by $(x,y)\mapsto(y,x)$ for $x,y\in F$. The latter case is known as the split case, and the unitary groups for hermitian spaces over these two quadratic étale algebras will be treated separately.
For $\epsilon\in\left\{\pm 1\right\}$, a finite dimensional $\epsilon$-hermitian space is a finite dimensional vector space $\mathrm{V}$ over $K$, equipped with an $\epsilon$-hermitian form, i.e., there exists a function $\left\langle\cdot,\cdot\right\rangle\colon\mathrm{V}\times\mathrm{V}\rightarrow K$, such that for all $w,x,y\in\mathrm{V}$, and $\alpha\in K$:
@@
If $\epsilon=1$, we call $\mathrm{V}$ hermitian, and if $\epsilon=-1$, we call $\mathrm{V}$ skew-hermitian. If $\delta\in K^{\times}$ is a trace zero element, that is, $\operatorname{tr}_{K/F}\left(\delta\right)=0$, consider the space $(\mathrm{W}^{\delta},\left\langle\cdot,\cdot\right\rangle_{\mathrm{W}^{\delta}})=(\mathrm{V},\left\langle\cdot,\cdot\right\rangle^{\delta})$ equipped with the form defined by $\left\langle x,y\right\rangle^{\delta}=\delta\left\langle x,y\right\rangle$. We have that the space $\mathrm{W}^{\delta}$ is a $-\epsilon$-hermitian space. Throughout the text, hermitian spaces will always be denoted using the symbol $\mathrm{V}$ or $\mathrm{H}$, and skew-hermitian spaces will always be denoted using the symbol $\mathrm{W}$.
We say that the space $\mathrm{V}$ is non-degenerate if for every $0\neq x\in\mathrm{V}$, there exists $y\in\mathrm{V}$, such that $\left\langle x,y\right\rangle\neq 0$. In this case, if $T\colon\mathrm{V}\rightarrow\mathrm{V}$ is a linear map, then there exists a unique linear map $T^{\ast}\colon\mathrm{V}\rightarrow\mathrm{V}$, such that
@@
If $\mathrm{V}$ is a non-degenerate one-dimensional $\epsilon$-hermitian space over $K$, we have that $\operatorname{U}(\mathrm{V})$ is isomorphic to the group
$K^{1}=\left\{x\in K^{\times}\mid xx^{c}=1\right\}$
by the map sending $x\in K^{1}$ to the multiplication by $x$ map $\mathrm{m}_{x}\colon\mathrm{V}\to\mathrm{V}$. We refer to the inverse of this map as the obvious isomorphism $\operatorname{U}(\mathrm{V})\to K^{1}$.
@@ 2.1.1. Invariants of $\epsilon$-hermitian spaces
Let $\mathrm{V}$ be a non-degenerate $\epsilon$-hermitian space such that $\dim_{K}\mathrm{V}=n$. The space $\mathrm{V}$ has a natural invariant known as the discriminant, which we explain below.
@@
$\det\left(\left\langle\cdot,\cdot\right\rangle\right)_{\mathbf{e}}\coloneq\det\left(\left\langle e_{i},e_{j}\right\rangle\right)_{ij}.$
If $\mathrm{V}$ is hermitian, then $\det\left(\left\langle\cdot,\cdot\right\rangle\right)_{\mathbf{e}}\in F^{\times}$ and it is well-known that the class of $\det(\left\langle\cdot,\cdot\right\rangle)_{\mathbf{e}}$ in $F^{\times}/N_{K/F}(K^{\times})$ is independent of the chosen basis $\mathbf{e}$. Hence, we may omit $\mathbf{e}$ from the notation and denote $\det\mathrm{V}=\det(\left\langle\cdot,\cdot\right\rangle)_{\mathbf{e}}\in F^{\times}/N_{K/F}\left(K^{\times}\right)$. Note that in the split case where $K=F\times F$, we have that $F^{\times}=N_{K/F}\left(K^{\times}\right)$, and therefore the invariant $\det\mathrm{V}$ is always trivial.
We define the discriminant of $\mathrm{V}$ as
@@
Suppose now that $F$ is a local field (we allow both archimedean and non-archimedean local fields) and that $K/F$ is a quadratic field extension. Using the non-trivial quadratic character $\omega_{K/F}$ of $F^{\times}/N_{K/F}(K^{\times})$ associated to the quadratic extension $K/F$ by local class field theory, we can encode the discriminant as a sign. For a hermitian space $\mathrm{V}$ as above we define
@@ 2.2. One-dimensional hermitian spaces with respect to an étale algebra
Let $L$ be an étale algebra of degree $n$ over $K$. We say that $\left(L,\sigma\right)$ is an étale algebra with involution if $\sigma\colon L\rightarrow L$ is an involution, such that for any $a\in K$, $\sigma(a)=a^{c}$. Given such an étale algebra with involution, we may define the notion of an $L$-hermitian (or $L$-skew-hermitian) space as in Section 2.1 by replacing $K$ with $L$ and the involution $x\mapsto x^{c}$ with the involution $\sigma$ in the definitions.
It is well-known that any étale algebra with involution is of the form $L=L_{E}=K\otimes_{F}E$, where $E$ is an étale algebra of degree $n$ over $F$ and $\sigma\colon L_{E}\rightarrow L_{E}$ is defined on pure tensors by
@@
for $k\in K$ and $h\in E$. Henceforth we will often write $\sigma\left(x\right)=x^{c}$ for $x\in L_{E}$.
Let $E$ be an étale algebra as above. For any $\lambda\in E^{\times}$, let $L_{E,\lambda}=(L_{E},\left\langle\cdot,\cdot\right\rangle_{L_{E,\lambda}})$ be the one-dimensional $L_{E}$-hermitian space equipped with the following form:
@@
As before, given a trace zero element $\delta\in L_{E}^{\times}$, we may define a one-dimensional $L_{E}$-skew-hermitian space $\mathrm{W}_{E,\lambda}^{\delta}=(L_{E},\left\langle\cdot,\cdot\right\rangle_{\mathrm{W}_{E,\lambda}^{\delta}})$ by setting $\left\langle x,y\right\rangle_{\mathrm{W}_{E,\lambda}^{\delta}}=\delta\left\langle x,y\right\rangle_{L_{E,\lambda}}$.
Consider the unitary group $\operatorname{U}(L_{E,\lambda})$, consisting of all $L_{E}$-linear maps that preserve the $L_{E}$-hermitian structure on $L_{E,\lambda}$. It is naturally isomorphic to the group of norm one elements of $L_{E}$:
@@
The isomorphism is given by the map sending $x\in L_{E}^{1}$ to the multiplication by $x$ map $\mathrm{m}_{x}\colon L_{E,\lambda}\to L_{E,\lambda}$. We refer to the inverse of this isomorphism as the obvious isomorphism $\operatorname{U}(L_{E,\lambda})\to L_{E}^{1}$. Similarly, the unitary group $\operatorname{U}(\mathrm{W}_{E,\lambda}^{\delta})$ is also naturally isomorphic to $L_{E}^{1}$, and we define the notion of the obvious isomorphism $\operatorname{U}(\mathrm{W}_{E,\lambda}^{\delta})\to L_{E}^{1}$. The group $L_{E}^{1}$ will be fundamental for defining maximal tori in unitary groups in the next sections.
If $E=\prod_{j=1}^{m}F_{j}$, where for every $j$, $F_{j}/F$ is a field extension, and $\lambda=\left(\lambda_{1},\dots,\lambda_{m}\right)\in\prod_{j=1}^{m}F_{j}^{\times}$, then we have that $\operatorname{U}(L_{E,\lambda})=\prod_{j=1}^{m}\operatorname{U}(L_{F_{j},\lambda_{j}})$, which is naturally isomorphic to $\prod_{j=1}^{m}L_{F_{j}}^{1}$. We also have a similar statement for the unitary group $\operatorname{U}(\mathrm{W}_{E,\lambda}^{\delta})$ of the skew-hermitian $L_{E}$-space $\mathrm{W}_{E,\lambda}^{\delta}$.
@@ 2.2.1. The discriminant of an étale algebra
$\det\left(\operatorname{tr}_{E/F}\left(\cdot\right)\right)_{\mathbf{e}}\coloneqq\det\left(\operatorname{tr}_{E/F}\left(e_{i}e_{j}\right)\right)_{i,j}.$
@@ 2.3.1. Classification of maximal tori of unitary groups
In this section, we utilize the space $\mathrm{V}_{E,\lambda}$ and the unitary group $\operatorname{U}(L_{E,\lambda})$ in order to classify the maximal tori of $\operatorname{U}(\mathrm{V})$ where $\mathrm{V}$ is a non-degenerate $n$-dimensional hermitian space.
For any $x\in L_{E}$, we may consider the $K$-linear map $\mathrm{m}_{x}\colon L_{E}\rightarrow L_{E}$ defined by $\mathrm{m}_{x}\left(y\right)=xy$. For any $\lambda\in E^{\times}$ and any $x\in L_{E}^{1}$, we have that the map $\mathrm{m}_{x}$ preserves the hermitian form $\left\langle\cdot,\cdot\right\rangle_{\lambda}$. Let
@@
By [27, Proposition 2.3] we have that $\mathrm{T}$ is a maximal torus in $\operatorname{U}(\mathrm{V}_{E,\lambda})$.
If $\mathrm{V}$ is a non-degenerate $n$-dimensional hermitian space, all maximal tori of $\operatorname{U}(\mathrm{V})$ can be described in a similar way. First, if $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$ is an isomorphism of hermitian spaces, then the map $\operatorname{U}(\mathrm{V}_{E,\lambda})\rightarrow\operatorname{U}(\mathrm{V})$ given by $g\mapsto r\circ g\circ r^{-1}$ is an isomorphism, and therefore
@@
is a maximal torus in $\operatorname{U}(\mathrm{V})$.
For the other direction, by [27, Proposition 2.3] if $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ is a maximal torus, then there exists an $n$-dimensional étale algebra $E$ and an embedding $i^{\prime}\colon L_{E}\rightarrow\operatorname{End}\left(\mathrm{V}\right)$, such that $\mathrm{T}=i^{\prime}\left(\operatorname{Res}_{L_{E}/F}\left(\operatorname{GL}_{1}\right)\right)\cap\operatorname{U}(\mathrm{V})$, or equivalently $\mathrm{T}=i^{\prime}\left(L_{E}^{1}\right)$. Finally, by [4, Proposition 1.4.1], if $E$ is an $n$-dimensional étale algebra over $F$, then there exists an embedding $i^{\prime}\colon L_{E}\rightarrow\operatorname{End}\left(\mathrm{V}\right)$ if and only if there exists $\lambda\in E^{\times}$ and an isomorphism of hermitian spaces $r\colon\mathrm{V}_{E,\lambda}\rightarrow\mathrm{V}$. By the proof of [4, Proposition 1.4.1], in this case we have $i^{\prime}\left(x\right)=r\circ\mathrm{m}_{x}\circ r^{-1}$ for any $x\in L_{E}$. We summarize these results in the following theorem.
@@
Hence, the statement in the split case is equivalent to the well-known fact that maximal tori in $\operatorname{GL}_{n}\left(F\right)$ are in bijection with subgroups of the form $\left\{\left[\mathrm{m}_{x}\right]_{\mathcal{B}_{E}}\mid x\in E^{\times}\right\}$ for some étale algebra $E$ of degree $n$ over $F$ and some $F$-basis $\mathcal{B}_{E}$ of $E$, where $\left[\mathrm{m}_{x}\right]_{\mathcal{B}_{E}}$ is the matrix representing $\mathrm{m}_{x}$ with respect to the basis $\mathcal{B}_{E}$. See for example [29, Proposition 3.2.29] or [31, Section 6.1].